OpenStudy (anonymous):

what is the solution to the nonhomogeneous equations: y''-9y=2xe^3x+5 y''+y=csc^2x

OpenStudy (amistre64):

this is more of a study site that guides you to get an answer rather than simply a site that will do your work for you.

OpenStudy (amistre64):

is that a system of equations , or are you simply trying to get 2 questions answered by one post?

OpenStudy (anonymous):

Its two questions in one post and I just want help specifically with finding the particular solution

OpenStudy (amistre64):

looks like they have constant coefficients; which means the homogenous part can be determined by a characteristic quadratic eqauation. the particular parts tho, hmm. the first appears to be workable from a table of "best guesses". 2xe^3x+5 best fits: yp = (ax + b)e^3x + c

OpenStudy (amistre64):

the second one tho would either have to take the long route or just go to a wonskian approach to shorten in up

OpenStudy (amistre64):

\[\begin{vmatrix}W_1&W&W_2\\ cos(x)&0&sin(x)\\-sin(x)&csc^2(x)&cos(x)\end{vmatrix}\]